OpenStudy (anonymous):

simplify sqrt (cos^2 x )+sqrt(x^3)

OpenStudy (anonymous):

\[\sqrt{\clubsuit^2}=|\clubsuit|\]

OpenStudy (anonymous):

\[\sqrt{\diamondsuit^3}=\diamondsuit\sqrt{\diamondsuit}\]

OpenStudy (anonymous):

?